# Long-Tailed Continual Learning For Visual Food Recognition

### v1

@@ Abstract
Deep learning based food recognition has achieved remarkable progress in predicting food types given an eating occasion image. However, there are two major obstacles that hinder deployment in real world scenario. First, as new foods appear sequentially overtime, a trained model needs to learn the new classes continuously without causing catastrophic forgetting for already learned knowledge of existing food types. Second, the distribution of food images in real life is usually long-tailed as a small number of popular food types are consumed more frequently than others, which can vary in different populations. This requires the food recognition method to learn from class-imbalanced data by improving the generalization ability on instance-rare food classes. In this work, we focus on long-tailed continual learning and aim to address both aforementioned challenges. As existing long-tailed food image datasets only consider healthy people population, we introduce two new benchmark food image datasets, VFN-INSULIN and VFN-T2D, which exhibits on the real world food consumption for insulin takers and individuals with type 2 diabetes without taking insulin, respectively. We propose a novel end-to-end framework for long-tailed continual learning, which effectively addresses the catastrophic forgetting by applying an additional predictor for knowledge distillation to avoid misalignment of representation during continual learning. We also introduce a novel data augmentation technique by integrating class-activation-map (CAM) and CutMix, which significantly improves the generalization ability for instance-rare food classes to address the class-imbalance issue. The proposed method is evaluated on Food101-LT, VFN-LT, VFN-INSULIN and VFN-T2D and show promising performance with large margin improvements compared with existing methods. We conduct an ablation study and discuss potential techniques that can further improve the performance, demonstrating great potential to deploy our method in real world food recognition applications.
@@ I Introduction
The emergence of modern deep learning technologies have enabled the automatic food nutrition content analysis, including image-based dietary assessment [1, 2, 3, 4], to help monitor and inform interventions to improve dietary intake and to prevent chronic diseases such as diabetes. As the first and fundamental step of image-based dietary assessment, food recognition aims to identify food types given an input image and the overall dietary assessment performance greatly relies on the precise food recognition results. Though existing deep learning based food recognition methods [5, 6, 7, 8, 9, 10] have achieved remarkable performance by training off-the-shelf Convolutional Neural Networks (e.g. ResNet[11]) using static datasets (e.g. Food-101[12], Food2K [13]), there are still two major challenges when applying them in a real world scenario including (i) how to update the model when new food classes appear sequentially overtime, and (ii) how to address the severe class-imbalance issue since the real life food images are usually in long-tailed distribution as shown in [14] where a minority of foods classes (i.e. instance-rich or head classes) are consumed more frequently than the remaining majority food classes (i.e. instance-rare or tail classes). The food recognition performance could drop dramatically without considering both two challenges. As shown in Figure 1, an ideal food recognition system in the real world should be able to learn new foods incrementally in long-tailed distribution with class-imbalance training samples at each incremental phase.
Figure 1: The overview of long-tailed continual learning. New classes with imbalanced distribution arrive sequentially over time at each incremental learning phase. The updated model should be able to learn new classes continuously and perform classification on all the classes seen so far.
@@
Continual learning, also known as incremental learning or lifelong learning, has been studied with the objective of learning new classes continuously without catastrophic forgetting [15] of learned knowledge. Compared with the naive solution of retraining the model from scratch whenever encountering a new class, continual learning is more practical as it only requires the data from new classes during the learning process, leading to the improvements in time, computation and memory efficiency [16]. However, the problem becomes more challenging if the data exhibits long-tailed distribution where the model needs to address both catastrophic forgetting and class-imbalance issues. Though the most recent work [17] aimed to solve this problem by introducing a 2-stage framework, the detached training process also presents challenges when implemented in real-world applications due to the inefficiency caused by manual fine-tuning of each individual stage. In addition, none of the existing work target on food images, which can be more challenging due to the intra-class variation and inter-class similarity.
@@
Existing continual learning methods show the effectiveness of applying knowledge distillation and storing a small fixed number of seen images as exemplars to mitigate catastrophic forgetting. However, both techniques become less effective in long-tailed distribution. Specifically, the knowledge distillation [18] may even harm the performance when teacher’s model is not trained on balanced data due to the bias in output logits as shown in a recent study [19]. On the other hand, distilling knowledge through learned representations impose a new challenge of feature space misalignment [20] as the learned representation needs to evolve during continual learning to accommodate new classes. Regarding using an exemplar set, most classes in long-tailed distribution may contain only a few training samples. Consequently, the overall performance may still be hindered even when all available samples are stored for instance-rare classes due to the poor generalization ability.
@@
In this work, we focus on designing an end-to-end long-tailed continual learning framework for visual food recognition. We leverage feature-based knowledge distillation while incorporating an additional prediction head that projects the current representation space to the past, which addresses the misalignment issue by providing more freedom to the student model and encourages the retention of the learned knowledge. In addition, inspired by the most recent work [21] that uses the context-rich information in head classes to help the tail classes, we introduce a new data augmentation technique by integrating class-activation-map (CAM) and CutMix [22], which cuts the most important region calculated by CAM in instance-rare classes data as foreground and pastes into the instance-rich classes images. With minimal computational overhead, this method significantly enhances the generalization capabilities of tail classes. We evaluate our method on existing long-tailed food image datasets including Food101-LT and VFN-LT. We further expand the VFN-LT to include two additional population groups, namely Insulin Takers and those with Type 2 diabetes without taking insulin, denoted as VFN-INSULIN and VFN-T2D, respectively. Our proposed framework achieves the best performance with a large improvement margin compared to existing methods while not requiring detached training stages. Finally, we conduct an ablation study to evaluate the effectiveness of each component in our proposed method and discuss potential techniques that can boost the accuracy for implementing in real world applications. The main contributions of this work are summarized as in the following:
  • We study the long-tailed continual learning in this work, which is closely related to food recognition in various real world scenarios. Despite its significance, this area of research is currently under-explored.
  • Two new benchmark long-tailed food image datasets including VFN-INSULIN and VFN-T2D are introduced, which exhibit the real world food consumption for the population of insulin takers and type 2 diabetes without taking insulin, respectively.
@@
  • We propose a novel framework to address both catastrophic forgetting and class-imbalance effectively by using feature-based knowledge distillation with a predication head and a novel CAM-based CutMix for data augmentation. In addition, an integrated loss is introduced to strike the balance between learning new classes and maintaining the learned knowledge.
@@
  • We conduct a set of comprehensive experiments on all long-tailed continual learning benchmarks for food recognition and discuss potential techniques that could boost the accuracy for facilitating the deployment in real world food recognition.
II Related Word
@@
In this section, we summarize existing methods that are most related to our work including the food recognition, long-tailed recognition and continual learning.
@@ II-A Food Recognition
Food image recognition is a challenging yet practical task that is related to various real world applications such as the image-based dietary assessment [23, 24, 25]. The performance of nutritional content analysis such as energy [26] or macronutrients are heavily reliant on the accuracy of recognizing and predicting the correct food types from food images. Most existing deep learning based work leverage off-the-shelf models [27, 28, 11, 29] and train on static food image datasets [12, 13, 30, 31, 32, 9]. In order to address the issue of inter-class similarity and intra-class variability, a manually built hierarchy based food recognition is proposed in [10]. Later, Mao et al[9] propose to construct a food hierarchy based on visual similarity and nutrition content [33] without requiring human efforts. In addition, food recognition has been studied under different scenarios such as ingredient recognition [34, 35], fine-grained recognition [36, 37, 38], few-shot learning [39, 40], long-tailed recognition [14, 41] and continual learning [42, 43]. However, none of the existing method are capable of learning new classes continuously in long-tailed distributions, which closely relates to real-world food recognition as new foods appear sequentially overtime with a minority of foods consumed more frequently than the others [14]. Though the most recent work [17] aims to integrate continual learning and long-tailed recognition, they decouple the training process to multi-stages and do not focus on food images. In this work, we target on long-tailed continual learning for visual food recognition to fill this gap and introduce a novel end-to-end framework to address both class-imbalance issue and catastrophic forgetting simultaneously.
@@ II-B Long-tailed Recognition
Image recognition in long-tailed distribution has been widely studied in recent years [44]. Existing work can be mainly categorized into two main groups including re-weighting and re-sampling based methods. The major challenge of long-tailed recognition is the imbalance training data between instance-rich (head) and instance-rare (tail) classes. The re-weighting based methods aim to balance the loss or gradients during the training process. The class-level re-weighting loss assigns weights to each class such as the Balanced Softmax loss [45], Label-Distribution-Aware Margin loss [46] where the inverse of class frequency [47] is widely used as the class weights. In addition, re-sampling based techniques aim to construct a balanced training set by over-sampling the tail classes or under-sampling the head classes. However, the naive over-sampling [48] by simply repeating the tail classes data for balanced training further intensifies the over-fitting issue and the naive under-sampling [49] by randomly removing the training data of head classes results in performance degradation. Therefore, most existing work performs efficient data augmentation techniques to improve the generalization ability of tail classes and achieve better overall recognition performance. Gao et al[41] propose Dynamic Mixup for multi-label long-tailed recognition problem, which dynamically adjusts the selection of images based on the previous training performance and set the label of synthetic image as the union of two images. The most recent work proposed CMO [21], which applies CutMix [22] for data augmentation by cutting the foreground region in tail classes images and paste in head classes background. The center idea of CMO is to leverage the context rich information from the head classes to help the generalization of tail classes. Later, He et al[14] improves the CMO to use visually similar image pairs and allows for multi-image CutMix to achieve improved performance. In spite of the efficiency of CutMix for data augmentation, one of the limitations is that it suffers from loss of semantic information of the original image since the cut region is generated randomly. Inspired by [50], we introduce a novel CAM based CutMix, which is able to seamlessly combine the images without losing semantic information, which will be illustrated in Section IV.
@@ II-C Continual Learning
Continual learning, also known as incremental learning or lifelong learning, has been studied under different scenarios such as class-incremental, task-incremental and domain-incremental learning as categorized in [51]. From the perspective of real world application, we primarily focus on class-incremental learning in this work, which aims to learn new classes continuously and perform classification on all classes seen so far during the inference phase, which does not require the task index or use multi-head classifier [52] compared to task-incremental learning. The domain-incremental learning instead does not involve new classes but only the domain shift for learned knowledge. The major challenge of class-incremental learning is catastrophic forgetting [15] where the updated model quickly forgets the learned knowledge due to the unavailability of learned classes data. There are two main groups of existing class-incremental learning methods including regularization based and replay based methods.
@@
Regularization based methods address forgetting by restricting the change of learned parameters during the learning of new classes. The initial work freeze the learned parameters in fully connected layer [53] or constraining the weights update [54], which also confines the model’s ability to learn from new data. Later, Li et al[55] proposed to use knowledge distillation [18] to maintain the learned knowledge by mimicking the output logits distribution from a teacher model, which is then improved in [56] by adding stronger constraints. In addition, feature-based knowledge distillation for class-incremental learning are applied in [57] by minimizing the discrepancy of learned representations between student and teacher models, which is further developed in [58] integrating the logits and feature-based distillation. However, existing knowledge distillation methods are developed on balanced data and becomes less efficient in long-tailed distribution. As revealed in [19], the knowledge distillation using output logits may even harm the overall performance if the teacher model is not trained on balanced data due to the bias towards instance-rich classes. Furthermore, directly applying feature based distillation may also impose new challenge such as feature space misalignment [20] due to the evolving of feature space when learning new classes especially in a long-tailed scenario where the data distribution may vary a lot for each incremental learning step. We address this problem by adding a prediction head to map the current representation space to the past, enabling more efficient knowledge transfer. Replay based methods assume the availability of a memory budget to store a small of number of learned classes data as exemplars to perform knowledge replay during the class-incremental learning. Herding algorithm [59] for exemplar selection based on the class mean vector is firstly proposed in [60], which becomes one of the most popular strategies in existing work [56, 61, 62, 57, 63]. Besides, He et al[42] further improves herding by applying an additional clustering step to select exemplars from each cluster based on herding, which mitigates the issue of intra-class variability. Despite the effectiveness of exemplar replay to maintain the learned knowledge, existing methods assume the training data is balanced and each class has more training data than the memory budget (e.g. 20 exemplars per class). However, the majority of classes in the long-tailed scenario only contain a few training samples, resulting in class-imbalance issues in the exemplar set and may even harm the overall performance when performing the knowledge replay. In this work, we address this issue by constructing a balanced exemplar set by augmenting the tail classes data with the proposed CAM based CutMix, which not only improves the efficiency of knowledge replay to maintain the knowledge but increases the generalization ability on tail classes to achieve better overall performance.
TABLE I: Example of matched food codes in VFN. (NFS denotes Not Further Specified.)
Food Type Food Code Main food description from FNDDS Additional food description
Apple 63101000 Apple, raw apple, NFS
Bagel 51180010 Bagel; flavored bagel; egg bagel; Bagel Thins
Baked potato 71100100 Potato, baked, NFS
Boiled egg 31103010 Egg, whole, boiled or poached soft boiled or hard boiled egg; egg, cooked, no fat or milk added
Chicken wings 24160110 Chicken wing, NFS as to cooking method drummette without sauce or seasoning
Cookies 53201000 Cookie, NFS
Fried rice 58150310 Rice, fried, NFS Chinese rice
Ice cream 13110000 Ice cream, NFS NFS as to flavor
Pancake 55100005 Pancakes, NFS hot cakes; flapjacks
Yogurt 11400000 Yogurt, NFS
III Long-tailed Food Datasets For Diabetes
The most recent work [14] introduced VFN-LT, which is a new long-tailed version of VFN [9] dataset for food recognition where the data distribution exhibits the real world food consumption frequencies [64, 65] for healthy people aged 18 to 65 in U.S.. However, around 34.2 million U.S. individuals (10.5 % of the U.S. population) have diabetes[66], which can cause various health problem such as heart disease, vision loss, and kidney disease. The management of diabetes is significantly influenced by diet, therefore, food recognition for predicting appropriate dietary choices to maintain control of diabetes is essential. Nevertheless, no existing studies have specifically focused on this special population so far. In this work, we aim to fill this gap by introducing two new benchmark long-tailed datasets including VFN-INSULIN and VFN-T2DM, which are constructed based on Viper FoodNet (VFN) datasets [9] and target on food image recognition for dietary assessment among insulin takers and those with type 2 diabetes without taking insulin, respectively.
The original VFN datasets consists of 74 most frequently consumed foods (exclude drink and beverages) selected based on What We Eat In America (WWEIA)  [^1] https://data.nal.usda.gov/dataset/what-we-eat-america-wweia-database. Following the similar process in [14], we first manually match each of the 74 food types in VFN [9] with one 8-digit USDA food codes from the Food and Nutrient Database for Dietary Studies (FNDDS) [^2] https://www.ars.usda.gov/northeast-area/beltsville-md-bhnrc/beltsville-human-nutrition-research-center/food-surveys-research-group/docs/fndds-download-databases/. Each 8-digit USDA food code represents a specific food item reported and consumed in the food supply . Table I shows examples of matched food codes from FNDDS in the VFN dataset. Next, food codes were labeled with their corresponding consumption frequency as determined by Lin et al [64] using the nationally representative dietary data collected through the National Health and Nutrition Examination Survey (NHANES)  [^3] https://www.cdc.gov/nchs/nhanes/index.html from 2009–2016 among U.S adults aged from 20 to 65. The consumption frequency represents how many times a particular food item was reported on one day in the particular age group among the U.S. population. As such, the frequency shows the prominence of the food in comparison with other foods reported as consumed in the population with 774 and 2,758 participants for insulin takers and type 2 diabetes without taking insulin. Finally, we construct VFN-INSULIN and VFN-T2D by reducing the number of training samples for food classes in original VFN based on the matched consumption frequency. Specifically, given the consumption frequency $f_{i}$ for the i-th food class, we calculate the number of training samples by (1)
$s_{i}=n_{i}\times\frac{f_{i}}{f_{max}}$ (1)
where $n_{i}$ is the original number of training samples in VFN for food class $i$ and $s_{i}$ denotes how many training images are kept for this class, which are randomly selected among original training data. $f_{max}$ refers to the maximum matched consumption frequency among the 74 food types in VFN datasets.
Overall, there are 2,056 training images from 74 food classes in VFN-INSULIN with a maximum of 288 training images per class and a minimum of 1 image per class. In addition, there are 2,145 training images in VFN-T2D with a maximum of 288 and a minimum 1 image per class. We observe that the food type Yeast bread, has a very high consumption among the represented adults and dominates the consumption frequency for both insulin takers and type 2 diabetes without taking insulin, resulting in the same imbalance ratio $\rho=\frac{max_{i}\{s_{i}\}}{min_{i}\{s_{i}\}}$ calculated as 288 for both datasets. However, the consumption frequencies of other food types may vary between the two groups. Figure 2 shows the distribution for each food type in VFN-INSULIN and VFN-T2D in descending order based on the number of training samples per class.

### v2

@@ Abstract
Deep learning-based food recognition has made significant progress in predicting food types from eating occasion images. However, two key challenges hinder real-world deployment: (1) continuously learning new food classes without forgetting previously learned ones, and (2) handling the long-tailed distribution of food images, where a few common classes and many more rare classes. To address these, food recognition methods should focus on long-tailed continual learning. In this work, We introduce a dataset that encompasses 186 American foods along with comprehensive annotations. We also introduce three new benchmark datasets, VFN186-LT, VFN186-INSULIN and VFN186-T2D, which reflect real-world food consumption for healthy populations, insulin takers and individuals with type 2 diabetes without taking insulin. We propose a novel end-to-end framework that improves the generalization ability for instance-rare food classes using a knowledge distillation-based predictor to avoid misalignment of representation during continual learning. Additionally, we introduce an augmentation technique by integrating class-activation-map (CAM) and CutMix to improve generalization on instance-rare food classes. Our method, evaluated on Food101-LT, VFN-LT, VFN186-LT, VFN186-INSULIN, and VFN186-T2DM, shows significant improvements over existing methods. An ablation study highlights further performance enhancements, demonstrating its potential for real-world food recognition applications.
@@ I Introduction
The emergence of modern deep learning technologies has enabled automatic food nutrition analysis, including image-based dietary assessment [1, 2, 3, 4], to monitor and improve dietary intake and prevent chronic diseases like diabetes. As the first step in this process, food recognition identifies food types from images, and accurate recognition is critical for overall assessment performance. Deep learning-based methods [5, 6, 7, 8] demonstrate remarkable performance by training off-the-shelf Convolutional Neural Networks (e.g. ResNet[9]) using static datasets (e.g. Food-101[10], Food2K [11]). However, two major challenges remain in real-world applications: (i) updating models as new food classes emerge over time, and (ii) addressing the severe class imbalance in long-tailed distributions, where a few classes (head classes) dominate consumption compared with most others (tail classes) [12, 13]. Failing to address these can significantly degrade performance.
@@
Continual learning, also known as incremental or lifelong learning, allows models to learn new classes continuously without catastrophic forgetting [14, 15, 16, 17, 18]. Unlike retraining from scratch whenever encountering a new class, continual learning is more practical, requiring only new class data, which improves time, computation, and memory efficiency [19]. The challenge intensifies when the data follows a long-tailed distribution [20, 21], requiring the model to address both catastrophic forgetting and class imbalance. While recent work [22] introduces a 2-stage framework to tackle this, its manual fine-tuning and detached training stages pose inefficiencies for real-world use. Additionally, existing methods have not been specifically applied to food images. This presents further challenges because food images often exhibit high intra-class variation and inter-class similarity, making it difficult to distinguish between different food items.
@@
Existing continual learning methods show the effectiveness of applying knowledge distillation and storing a small fixed number of seen images as exemplars to mitigate catastrophic forgetting. However, both techniques become less effective in the long-tailed distribution. Knowledge distillation [23],[24] may even harm the performance when the teacher’s model is not trained on balanced data due to the bias in output logits as shown in a recent study [25]. On the other hand, distilling knowledge through learned representations imposes a new challenge of feature space misalignment [26] as the learned representation needs to evolve during continual learning to accommodate new classes. Regarding using an exemplar set, most classes in long-tailed distribution may contain only a few training samples. Consequently, the overall performance may still be hindered even when all available samples are stored for instance-rare classes due to the poor generalization ability.
@@
In this work, we focus on designing an end-to-end long-tailed continual learning framework for visual food recognition. We leverage feature-based knowledge distillation while incorporating an additional prediction head that projects the current representation space to the past. This addresses the misalignment issue by providing more freedom to the student model and encourages the retention of the learned knowledge. In addition, inspired by the most recent work [27] that uses the context-rich information in head classes to help the tail classes, we introduce a new data augmentation technique by integrating class-activation-map (CAM) and CutMix [28], which cuts the most important region calculated by CAM in instance-rare classes data as foreground and pastes into the instance-rich classes images. With minimal computational overhead, this method significantly enhances the generalization capabilities of tail classes. We evaluate our method on existing long-tailed food image datasets including Food101-LT and VFN-LT [12]. Additionally, we developed VFN186 based on the original VFN [7], expanding the initial 74 food categories by adding 112 more. This allows for a more comprehensive coverage of the typical American diet [29]. Furthermore, we derive three long-tailed versions of VFN186, referred to as VFN186-LT, VFN186-INSULIN, and VFN186-T2D, based on different population groups, namely healthy populations, Insulin Takers, and those with Type 2 diabetes without taking insulin. Our proposed framework achieves the best performance with a large improvement margin compared to existing methods while not requiring detached training stages. Finally, we conduct an ablation study to evaluate the effectiveness of each component in our method and discuss potential techniques that can boost the accuracy for real-world applications. The main contributions of this work are summarized in the following:
  • We introduce the VFN186 food dataset, which contains 186 most frequently consumed food types in America. Additionally, we introduce three new long-tailed benchmark datasets, which reflect the food consumption patterns of different populations. The dataset will be public.
@@
  • We propose a novel framework that utilizes feature-based knowledge distillation with a prediction head and a novel CAM-based CutMix for data augmentation, and an integrated loss function to address catastrophic forgetting and class imbalance.
@@
  • We conduct extensive experiments on all long-tailed continual learning benchmarks for food recognition and discuss potential techniques to enhance accuracy that could boost the accuracy for facilitating the deployment in real-world food recognition.
II Related Work
@@
In this section, we summarize existing methods most related to our work including food recognition, long-tailed recognition, and continual learning.
@@ II-A Food Recognition
Food image recognition is a challenging yet practical task with applications like image-based dietary assessment [30, 31], where accurate recognition is crucial for nutritional content analysis, such as energy and macronutrients [32, 33, 34, 35]. Most existing deep learning based work leverage off-the-shelf models [36, 37, 9, 38] and train on static food image datasets [10, 11, 39, 40, 41, 7]. To address inter-class similarity and intra-class variability, various hierarchy-based approaches have been proposed [8, 7, 42]. While food recognition has been studied in scenarios like ingredient recognition [43], fine-grained recognition [44, 45], few-shot learning [46], long-tailed recognition [12, 47, 13], and continual learning [48, 20], no existing methods continuously learn new classes in long-tailed distributions, which is critical for real-world applications [12]. Recent work [22] attempted to integrate continual learning with long-tailed recognition but used a multi-stage training process and did not focus on food images. In this work, we target long-tailed continual learning for visual food recognition, introducing a novel end-to-end framework to address both class imbalance and catastrophic forgetting simultaneously.
@@ II-B Long-tailed Recognition
Existing work on image recognition in long-tailed distributions can be categorized into two main groups: re-weighting and re-sampling. The major challenge is the imbalance between instance-rich (head) and instance-rare (tail) classes [49]. Re-weighting methods balance the loss or gradients during training, with a class-level re-weighting loss like Balanced Softmax [50] and Label-Distribution-Aware Margin loss [51]. In addition, re-sampling based techniques construct a balanced training set by over-sampling tail classes or under-sampling head classes, but naive over-sampling [52] and under-sampling [53] can lead to overfitting or performance degradation. Therefore, most existing work performs data augmentation to improve the generalization ability of tail classes and achieve better overall performance. Gao et al[47] propose Dynamic Mixup for multi-label long-tailed recognition problem, which dynamically adjusts the selection of images based on the previous training performance and set the label of synthetic image as the union of two images. CMO [27] applies CutMix [28] for data augmentation by cutting the foreground region in tail classes images and pasting in head classes background. The center idea of CMO is to leverage the context rich information from the head classes to help the generalization of tail classes. Later, He et al[12] improves the CMO to use visually similar image pairs and allows for multi-image CutMix to achieve improved performance. In spite of the efficiency of CutMix for data augmentation, one of the limitations is that it suffers from loss of semantic information of the original image since the cut region is generated randomly. Inspired by [54], we introduce a novel CAM-based CutMix, which combines the images seamlessly without losing semantic information, as detailed in Section IV.
@@ II-C Continual Learning
Continual learning, also known as incremental or lifelong learning, has been explored in scenarios like class-incremental, task-incremental, and domain-incremental learning [55]. This work focuses on class-incremental learning, which is key for real-world applications. It involves continuously learning new classes and classifying all previously seen classes during inference, without using task indexes or multi-head classifiers as in task-incremental learning [56]. Unlike domain-incremental learning, which handles domain shifts without new classes, class-incremental learning faces the challenge of catastrophic forgetting [14], where the model forgets previous knowledge due to the lack of data from learned classes [57]. To address this, existing methods are mainly divided into regularization-based and replay-based approaches.
@@
Regularization-based methods address forgetting by limiting changes to learned parameters while learning new classes. Initial work froze or constrained parameter updates [58, 59], limiting the model’s ability to learn new data. Later, Li et al[60] used knowledge distillation[23] to preserve learned knowledge by mimicking the teacher model’s output logits. Feature-based distillation minimized representation discrepancies [61] of learned representations between student and teacher models, which is further developed in [62] integrating the logits and feature-based distillation. However, the knowledge distillation using output logits may even harm the overall performance if the teacher model is not trained on balanced data due to the bias towards instance-rich classes. Furthermore, direct feature-based distillation also faces challenges like feature space misalignment [26] due to the evolving of feature space when learning new classes especially in a long-tailed scenario where the data distribution may vary a lot for each incremental learning step. We address this problem by adding a prediction head to map the current representation space to the past, enabling more efficient knowledge transfer.
Replay-based methods use a memory buffer to store exemplar data for knowledge replay during class-incremental learning. The herding algorithm [63] selects exemplars based on class mean vectors and is widely used [64, 65, 66, 67]. However, these methods assume balanced training data and sufficient samples per class compared with the memory budget (e.g. 20 exemplars per class), which isn’t the case in long-tailed scenarios. It can lead to class imbalance in the exemplar set and harm overall performance. We address this issue by constructing a balanced exemplar set by augmenting the tail class data with the proposed CAM-based CutMix, which augments tail class data, improving knowledge replay efficiency and generalization on tail classes for better overall performance.
II-D Continual Learning with Pre-trained Models
Leveraging large-scale pre-trained vision transformers for continual learning has emerged as a promising direction. Recent works optimize prompt parameters, a small set of learnable weights, to guide model predictions without storing past examples, which enable efficient knowledge encoding and transfer across tasks. L2P [68] pioneers prompt-based continual learning, which encodes knowledge through learnable prompt parameters to facilitate efficient adaptation. Wang et al proposes DualPrompt [69], which learn two sets of disjoint prompt spaces to encode task-invariant and task-specific instructions respectively. CODA-Prompt [70] further refines them using decomposed prompts, which dynamically assemble into attention-conditioned prompts optimized in an end-to-end manner. These methods leverage pre-trained Vision Transformer (ViT) and demonstrate strong performance on benchmark datasets, highlighting the potential of prompt-based approaches in continual learning [70].
III Datasets
III-A VFN186 Dataset
To encompass a broader range of food types, we expanded VFN [7] to include 186 food categories [^1] The dataset is available at https://github.com/JiangpengHe/VFN186. Similar to VFN, we select an additional 112 commonly consumed foods by Americans based on What We Eat In America (WWEIA) [29] database. This expansion makes the dataset more comprehensive and robust for training practical models with broader applicability. Specifically, similar to [12, 71], we first match each of the 186 food types in VFN186 with one 8-digit USDA food code from the Food and Nutrient Database for Dietary Studies (FNDDS) where each 8-digit USDA food code represents a specific food item in the food supply. Then we use a semi-automatic data collection system to crawl specific types of food images from the Google Image website based on food labels. Next, we employ a trained Faster R-CNN [72] to remove noisy images. The remaining images were processed using an online crowdsourcing tool, where food items were boxed and labeled with their corresponding categories. Through this process, we expand the VFN dataset and created the VFN186 dataset, which includes 186 food types and 70230 images.
III-B Long-tailed Food Datasets For Different Populations
While our VFN186 dataset provides rich value for various downstream tasks, in this work, we primarily focus on its long-tailed version for continual learning, which leverages its strength of matched food codes from the nutrition database and addresses the challenges of real-world food data distribution across different populations. Specifically, we generated three long-tailed versions of our VFN186 dataset. First, VFN186-LT follows the methodology of [12], reflecting real-world food consumption frequencies among healthy individuals aged 18 to 65 in U.S. as reported by [73].
Additionally, we developed VFN186-INSULIN and VFN186-T2D, designed for dietary assessment among insulin takers and those with type 2 diabetes without taking insulin, respectively. The motivation of developing VFN186-INSULIN and VFN186-T2D lies in addressing critical gap in food recognition research for the approximately 34.2 million U.S. individuals (10.5% of the population) with diabetes [74, 75]. Given the crucial role of diet in managing diabetes and its associated health complications, these population-specific long-tailed datasets aim to enhance the practical applicability of food recognition models in real-world scenarios.